OpenStudy (anonymous):

HELP!!! Solve by completing the square. x² - 2x - 24 = 0

OpenStudy (anonymous):

x² - 2x - 24 = 0 x² - 2x+1-1-24=0 (adding and subtracting 1) \[(x-1)^{2}-25=0\] \[(x-1)^{2}=25\] \[(x-1)^{2}=5^{2}\] x-1=5 x=6

OpenStudy (anonymous):

also x-1= -5 x=-4

OpenStudy (anonymous):

What does that make the aswer

OpenStudy (anonymous):

when we remove square root absolute(mod) function come as 3\[3^{2}=9 also (-3)^{2}=9\]